# Functional vs. Object-Oriented: Comparing How Programming Paradigms Affect the Architectural Characteristics of Systems

### v1

@@ 7. Qualitative Analysis
The qualitative analysis compares functionally equivalent code snippets from the Kotlin and Scala proof of concepts, evaluating them according to the architectural characteristics defined in table˜1. This methodology is self-ethnographic, since it presents the assessments of the authors of the paper, who developed the proofs of concept. Because of space constraints, this sections focuses on the architectural characteristics described in section˜2 and section˜2.
7.1. Extensibility
@@
In Kotlin, the Wallet entity is structured using OOP principles like inheritance and polymorphism. The abstract class Wallet serves as a blueprint for all specific wallet types: Real Money Wallet, Investment Wallet, and Emergency Funds Wallet. This is illustrated in section˜7.1
In Kotlin, getAvailableBalance is an abstract method of the Wallet class, which encapsulates the logic for querying balances from the ledger. All subclasses must implement their specific logic for computing balances. This ensures consistency across wallet types while allowing each subclass to customize its behavior. Such approach leverages polymorphism: the behavior of a method depends on the subclass of the object being called.
@@
In Scala, the Wallet entity is structured using FP principles like immutable data structures. The case class Wallet consolidates all possible attributes, and contains an additional field WalletType that distinguishes the wallet derivations within the system. This is illustrated in section˜7.1.
@@
In Scala, getAvailableBalance is a standalone function in the WalletService class, which encapsulates the logic for querying balances from the ledger. For every variant of the WalletType enumeration, the function specifies a corresponding list of ledger queries to compute the available balance. This provides behavior tailored to each wallet type without requiring separate case classes. Such approach leverages pattern matching: the walletType field maps to different balance computation logic.
@@ 7.2. Reusability
In Kotlin, the implementation of the retry logic is built around a while loop. This is illustrated in section˜7.2. It leverages the counter variable attempts to track the number of retries for each transaction and the flag variable success to indicate whether the transaction processing succeeded.
@@
In Scala, the implementation of the retry logic is build around a higher-order function named retry. This is illustrated in section˜7.2. It takes a function f: => Either[A, B] and an integer n as parameters to execute f up to n times, early returning Right if any attempt succeeds, or returning Left if all n attempts fail.
@@
7.3. Error Handling and Propagation
@@
In Kotlin, the hold Transaction is created through the processTransaction method, which can throw exceptions. These exceptions must be explicitly caught in the invest method to map them to a corresponding WalletService exception. Additionally, within the catch blocks, it is necessary to ensure the Transaction transitions to the appropriate state. This is illustrated in section˜7.3.
However, if a new exception type is created in the future, developers must explicitly update the try-catch block adding a new catch case, otherwise risking unhandled runtime exceptions if they forget about it.
Additionally, as a try-catch block might or not handle all the error cases, reasoning about the program becomes more challenging since the code does not clearly express the potential failure paths.
@@
In Scala, errors are emitted with via the monadic type Either. This approach allows them to be explicitly modeled as part of the method’s return type, making it clear to developers that a method can fail. This is illustrated in section˜7.3.
The use of Either enhances type safety by enforcing that the caller must handle both success and failure cases, ensuring that no error is missed during development.
The type-safe nature of Either eliminates the risk of unhandled runtime errors, promoting more predictable and reliable error handling within the system.
@@ 8. Quantitative Analysis
This section was designed to measure the developer’s impressions about the two Digital Wallet system proof of concepts. Each question includes a brief explanation of a functionality from the Digital Wallet system, followed by two code snippets: the first from the Kotlin PoC, and the second from the Scala PoC. For each code snippet, participants were asked to evaluate the six architectural characteristics summarized in table˜1. They answered the question: “From the code, how easy it is to…” using a Likert scale with five steps, from Very Easy to Very Hard.
At the publication of the bachelor thesis which inspired this paper, the survey received eight responses. Considering space constraints, this section will only address two of four questions proposed in the questionnaire, which intersect with some of the code previously analyzed in section˜7.
@@ 8.1. Question #1 — retryBatch
Question #1 collected feedback on the implementation of the retryBatch, whose code was previously explored in section˜7.2.
[…]
Kotlin vs. Scala — retryBatch Refer to caption Refer to caption
Figure 6. Kotlin vs. Scala — retryBatch
The results of Question #1 reveal a notable differences in the architectural characteristics of error handling, error propagation, and readability.
@@ 8.2. Question #3 — invest
Question #3 collected feedback on the implementation of the invest logic, whose code was previously explored in section˜7.3.
[…]
Kotlin vs. Scala — invest Refer to caption Refer to caption
Figure 7. Kotlin vs. Scala — invest
The results of Question #3 reveal that extensibility is slightly favored in the Scala PoC compared to the Kotlin PoC. On the other hand, testability showed a significant advantage in favor of Kotlin.
@@ 9. Discussion
With the findings derived from both the qualitative and quantitative analyses, it is possible to finally answer the research subquestions proposed in section˜2.
RQ1.1: How do functional and object-oriented paradigms impact the extensibility of a system?
@@
RQ1.2: How do functional and object-oriented paradigms impact the reusability of a system?
@@
RQ1.3: How do the functional and object-oriented paradigms impact error handling / propagation of a system?
@@ 10. Threats to Validity
Assessing threats to validity is essential to acknowledge the limitations that influence any research.
@@
The evaluation of programming paradigms in this research may be influenced by the choice of programming languages representing each paradigm. Kotlin and Scala were chosen as modern, multi-platform languages that run on the Java Virtual Machine (JVM), ensuring a level of syntactic compatibility between the programming languages.
Despite Kotlin supporting functional features and Scala supporting object-oriented features, the research intentionally limited Kotlin and Scala to their selected domains.
@@ 11. Conclusion
This paper introduced the Digital Wallet system, implemented twice: once in Kotlin (OOP), once in Scala (FP). These proofs of concept were analyzed in a self-ethnographic qualitative analysis and a survey-based quantitative analysis, to understand how the object-oriented and functional programming paradigms impact six architectural characteristics presented in table˜1.
For future works, this research can be extended to target other architectural characteristics, such as scalability and security. Additionally, the current results can be enhanced by expanding the survey, thus addressing some of the known threats to validity.

### v3

@@ 8. Qualitative Analysis
The qualitative analysis compares functionally equivalent code snippets from the Kotlin and Scala proofs of concept, evaluating them according to the architectural characteristics focused on by the research sub-questions. This methodology is self-ethnographic, since it presents the assessments of the authors who led the development process.
8.1. Extensibility [section 1]
@@
List of codes 1 Wallet class inheritance hierarchy in Kotlin :
In Kotlin, the Wallet entity is structured using OOP principles like inheritance and polymorphism. The abstract class Wallet serves as a blueprint for all specific wallet types: Real Money Wallet, Investment Wallet, and Emergency Funds Wallet. This is illustrated in listing 1.
In Kotlin, getAvailableBalance is an abstract method of the Wallet class, which encapsulates the logic for querying balances from the ledger. All subclasses must implement their specific logic for computing balances. This ensures consistency across wallet types while allowing each subclass to customize its behavior. Such an approach leverages polymorphism: the behavior of a method depends on the subclass of the object being called.
@@
In Scala, the Wallet entity is structured using FP principles like immutable data structures. The case class Wallet consolidates all possible attributes and contains an additional field WalletType that distinguishes the wallet derivations within the system. This is illustrated in listing 2.
@@
List of codes 2 Wallet type and factory pure function in Scala :
In Scala, getAvailableBalance is a standalone function in the WalletService class, which encapsulates the logic for querying balances from the ledger. For every variant of the WalletType enumeration, the function specifies a corresponding list of ledger queries to compute the available balance. This provides behavior tailored to each wallet type without requiring separate case classes. Such an approach leverages pattern matching: the walletType field maps to different balance computation logic.
@@ 8.2. Reusability [section 1]
List of codes 3 retryBatch built with a for loop in Kotlin
In Kotlin, the implementation of the retry logic is built around a while loop. This is illustrated in listing 3. It leverages the counter variable attempts to track the number of retries for each transaction and the flag variable success to indicate whether the transaction processing succeeded.
@@
List of codes 4 retryBatch built over the HOF retry in Scala
In Scala, the implementation of the retry logic is built around a higher-order function named retry. This is illustrated in listing 4. It takes a function f: => Either[A, B] and an integer n as parameters to execute f up to n times, early returning Right if any attempt succeeds, or returning Left if all n attempts fail.
@@
8.3. Error Handling and Propagation [section 1]
@@
List of codes 5 invest method using exceptions in Kotlin
In Kotlin, the hold Transaction is created through the processTransaction method, which can throw exceptions. These exceptions must be explicitly caught in the invest method to map them to a corresponding WalletService exception. Additionally, within the catch blocks, it is necessary to ensure the Transaction transitions to the appropriate state. This is illustrated in listing 5.
However, if a new exception type is created in the future, developers must explicitly update the try-catch block, adding a new catch case, otherwise risking unhandled runtime exceptions if they forget about it.
Additionally, as a try-catch block might not handle all the error cases, reasoning about the program becomes more challenging since the code does not clearly express the potential failure paths.
@@
List of codes 6 invest function using monadic types in Scala
In Scala, errors are emitted via the monadic type Either. This approach allows them to be explicitly modeled as part of the method’s return type, making it clear to developers that a method can fail. This is illustrated in listing 6.
The use of Either enhances type safety by enforcing that the caller must handle both success and failure cases, ensuring that no error is missed during development. This promotes more predictable and reliable error handling within the system.
@@ 9. Quantitative Analysis
This section was designed to measure the developer’s impressions about the two Digital Wallet system proof of concepts. Each question includes a brief explanation of a functionality from the Digital Wallet system, followed by two code snippets: the first from the Kotlin PoC and the second from the Scala PoC. For each code snippet, participants were asked to evaluate the six architectural characteristics summarized in table 1. They answered the question: “From the code, how easy is it to…” using a Likert scale with five steps, from Very Easy to Very Hard.
At the publication of the bachelor thesis that inspired this paper (Dias et al., 2024), the survey received eight responses. Considering space constraints, this section will only address two of four questions proposed in the questionnaire, which intersect with some of the code previously analyzed in section 8. For a full analysis, please refer to the bachelor thesis (Dias et al., 2024).
@@ 9.1. Question #1 — retryBatch
Question #1 collected feedback on the implementation of the retryBatch, whose code was explored in listings 3 and 4.
The results of Question #1 are shown in figure 5. They reveal a notable difference in the architectural characteristics of error handling, error propagation, and readability.
@@ 9.2. Question #3 — invest
Question #3 collected feedback on the implementation of the invest logic, whose code was explored in listings 5 and 6.
The results of the Question #3 are shown in figure 6. They reveal that extensibility is slightly favored in the Scala PoC compared to the Kotlin PoC. On the other hand, testability showed a significant advantage in favor of Kotlin.
@@ 10. Discussion
Based on the findings from the qualitative and quantitative analyses, this section discusses the research subquestions.
sRQ1: How do the functional and object-oriented paradigms impact the extensibility of a system?
@@
sRQ2: How do the functional and object-oriented paradigms impact the reusability of a system?
@@
sRQ3: How do the functional and object-oriented paradigms impact error handling and propagation of a system?
@@ 11. Threats to Validity
The threats to validity acknowledge the limitations of this research.
@@
The evaluation of programming paradigms in this research may be influenced by the choice of programming languages representing each paradigm. Kotlin and Scala were chosen as modern, multi-platform languages that run on the Java Virtual Machine (JVM), ensuring a level of syntactic compatibility between the programming languages. Despite Kotlin supporting FP features and Scala supporting OOP features, this research intentionally limited both to their selected domains.
@@ 12. Conclusion
RQ: How do the functional and object-oriented paradigms impact different architectural characteristics of a system?
This paper compared the impact of adopting OOP versus FP over six architectural characteristics of software systems. It introduced the Digital Wallet system, implemented twice: once in Kotlin (OOP), once in Scala (FP). The use of these modern, popular programming languages for the research will hopefully reach the interest of practitioners, who may understand the results and then be better informed when choosing a certain paradigm.
For future work, this research can be extended to target other architectural characteristics, such as scalability and security. Additionally, the current results can be enhanced by expanding the survey, thus addressing some of the threats to validity.